Write the equation of the line passing through each of the given pairs of points. Write your result in slope-intercept form, where possible.
(20) (-1,3) and (4, -2)
(24) (2, -3) and (2,4)

Answer by venugopalramana(3286)

Write the equation of the line passing through each of the given pairs of points. Write your result in slope-intercept form, where possible.
STD. EQN. OF LINE THROUGH P(X1,Y1) AND Q(X2,Y2) IS
(Y-Y1)/(Y2-Y1)=(X-X1)/(X2-X1)
(20) (-1,3) and (4, -2)
(Y-3)/(-2-3)=(X+1)/(4+1)
(Y-3)/-5=(X+1)/5
Y-3=-5(X+1)/5=-X-1
Y=-X+2
(24) (2, -3) and (2,4)
(Y+3)/(4+3)=(X-2)/(2-2)
(Y+3)/7=(X-2)/0
X-2=0